# An Optimal Framework for Constructing Lie-Algebra Generator Pools: Application to Variational Quantum Eigensolvers for Chemistry

### v1

@@ I Introduction
The evolution of quantum states is fundamentally described by unitary transformations that form a Lie group. The generators of these evolutions are elements of the corresponding Lie Algebra, which can be understood as the tangent space at the identity element of the Lie group when seen as a smooth manifold. Lie Algebras are central in quantum computation, where any transformation can be realized as a sequence of unitary evolutions generated by operators from a specific Lie Algebra, typically $\mathfrak{su}(2^{N})$ for an $N$-qubit system. A key research area involves understanding the generation of these algebras from a minimal set of operators. For instance, when using a set of products of Pauli operators, it is known that a minimum of $2N+1$ elements in this set are required to generate the full basis of $\mathfrak{su}(2^{N})$ through nested commutation relations. The span of this basis is known as Dynamical Lie Algebra (DLA), a term used in the context of quantum optimal control [Albertini, Wiersema2024, Goh2025, fuclass, Smith2025, class2, Kazi2025], and later adopted in quantum computing. Consequently, significant past work has aimed to investigate specific properties and applications of these bases. Broadly, these efforts can be categorized into two themes: foundational classification [Wiersema2024, Goh2025, fuclass, Smith2025, class2, Kazi2025] and practical application in variational quantum algorithms [Shkolnikov2023, Sapova2022, Haidar2025].
@@ II.1 Summary of Main Lie Algebra Results
The following table summarizes different approaches from the literature for verifying MCPs, alongside our proposed method. For more details on these existing methods, we refer the reader to Sup.Mat. LABEL:supp-poolconstr:
@@ II.2 Preliminaries
The proof of this proposition is given in Sup. Mat.  LABEL:supp-prop_com.
@@ II.3 Main theorems
For sake of clarity and brevity, we will present our results without proofs. Please refer to Sup. Mat. LABEL:supp-Prem for additional details and fully worked-out proofs.
@@
The proofs for Propositions 13 and 15 are detailed in Sup. M. LABEL:supp-main. The idea behind these propositions is that we can find a sequence of contractions, represented by a matrix $\mathbf{P}$, that transforms $\mathbf{\Gamma}_{\mathcal{A}}$ into an equivalent matrix $\mathbf{\Gamma}_{equiv}$ with a known classification.
@@
The corresponding proof is given in Sup. Mat. LABEL:supp-main.
@@
The proofs of these theorems are given in Sup.Mat. LABEL:supp-main.
@@ II.4 Outlook on Applications: Quantum Chemistry on Quantum Computers
where $M\in\mathbb{N}^{*}$, and $\boldsymbol{\theta}=\left\{\theta_{i}\right\}_{i=1}^{M},~\theta_{i}\in\mathbb{R}$ is a list of parameters that completely characterize the state. $U_{i},i\in[M]$ are unitary operators, can be defined also as excitation operators acting on a reference state $|\Psi_{0}\rangle$. This algorithm’s distinct feature is its iterative approach to constructing this parametrized quantum state. At each iteration, the state is expanded by appending an exponentiated generator, which is carefully selected from a set of operator pools. The general workflow of this algorithm is described in Sup. Mat. LABEL:supp-Adapt. There are many types of operator pools that can be used to construct the ADAPT-VQE ansatz, however one significant measurement overhead of ADAPT-VQE is the need to measure the energy gradient for every operator in the predefined operator pool which traditionally scales $\mathcal{O}(N^{4})$ yielding a measurement cost of $\mathcal{O}(N^{8})$. The work of [Shkolnikov2023], introduces a method to mitigate this cost by carefully constructing the operator pool, such that they use:
@@
Symmetries and properties of the pool: For molecular simulations with real-valued ansatz, the operator pool generators are restricted to odd Pauli strings (i.e odd number of Y operators in the string). The complete set of these generators forms the $\mathfrak{so}(2^{N})$ Lie Algebra, see Sup. Mat LABEL:supp-properties. However, molecular Hamiltonians possess additional inherent symmetries, such as the conservation of particle number, total spin, spin projection, and spatial point-group symmetries, see Sup. Mat. LABEL:supp-properties for more details. If the pool generators do not respect these symmetries, gradient-based algorithms like ADAPT-VQE can fail. This failure manifests as vanishing gradients, symmetry roadblock [Shkolnikov2023], preventing the algorithm from starting or converging.
@@
It is worth mentioning that the set of operators that preserves all these molecular symmetries is not necessarily closed under the Lie brackets, i.e the commutator of two symmetric operators may produce a new operator that violates the desired symmetry (e.g. particle number Sup. Mat. LABEL:supp-properties). Given that completeness requires the final pool to generate a mathematically closed Lie Algebra, the algebra generated from these starters will unavoidably expand to include operators that do not preserve all of the initial symmetries. We have therefore investigated which of these properties form a closed subalgebra. Our analysis indicates that the largest subalgebra that respects these constraints is defined by the two properties: the odd string requirement and the even flip property, the proof is detailed in Sup. Mat. LABEL:supp-properties. We found numerically that the MCPs corresponding to this specific Lie subalgebra, which is a subalgebra of $\mathfrak{so}(2^{N})$, have $\mathbf{\Gamma}_{\mathcal{A}}$ matrices of rank 2N42N-4. This finding is currently numerical, as a formal proof and detailed characterization of this Lie subalgebra are beyond the scope of this work.
@@ II.5 Numerical results
These methods are performed across systems ranging from 12 to 26 qubits, from weak to strongly correlated systems. The specific molecular benchmarks include: LiH ($R=1.3\text{{\AA }}$, 12 qubits), the hydrogen chain H6 ($R=3.0\text{\AA }$, 12 qubits), H8 ($R=0.8\text{\AA }$, 16 qubits), and H2O (26 qubits, geometry defined in Sup. Mat. LABEL:supp-geo). All simulations were performed using our Hyperion statevector emulator [hyperion]. We employed the minimal basis set (sto-3g) for most systems, while the 66-$31g$ basis set was used specifically for the H2O molecule. We assess ansatz performance across iterations using four metrics: energy accuracy, CNOT count, function evaluations (including the gradient measurements for ADAPT-VQE approaches), and the number of variational parameters.
@@
Figure 2: Convergence plots for ADAPT-VQE and MB-ADAPT-VQE methods across various molecules: LiH ($R=1.3\text{{\AA }}$, 12 qubits), the hydrogen chain H6 ($R=3.0\text{\AA }$, 12 qubits), H8 ($R=0.8\text{\AA }$, 16 qubits), and H2O (26 qubits, geometry defined in Sup. Mat. LABEL:supp-geo). The ADAPT-VQE method uses a QEB pool of size $\mathcal{O}(N^{4})$. The composite MB-ADAPT-VQE method employs batched-ADAPT-VQE with efficient construction of MCPs of size $\mathcal{O}(N)$ for various values of $k$. Plots display energy errors versus CNOT count, function evaluations, and parameter count for different molecules.
@@
Figure 3: Convergence plots for NI-DUCC-VQE method across various molecules using different pool operators: the hydrogen chain H6 ($R=3.0\text{\AA }$, 12 qubits), H8 ($R=0.8\text{\AA }$, 16 qubits), and H2O (26 qubits, geometry defined in Sup. Mat. LABEL:supp-geo). Plots display energy errors versus function evaluations.
@@ IV Methods
We give in Figure 5, an illustrated scheme on different steps towards a user-defined complete pool. As mentioned above, for our user-defined pool, which targets molecular Hamiltonians, the process begins with an initial set of starters. These starters must preserve specific symmetries and properties (see Sup. Mat. LABEL:supp-properties), such as odd strings, even flip, particle number, and spin. These starters can be selected in several ways: random selection from subgroups $G\subseteq i\mathcal{P}_{N}^{*}$ that preserve desired symmetries, random selection from UCCSD operator pool ( accounting only for single and double excitations) [Sapova2022, Haidar2025], or using the UCCSD pool combined with prescreening techniques to select the most significant contributors to energy convergence [Haidar2025]. Once the starters are chosen, the complete pool is built iteratively. We add new elements that preserve the required symmetries, provided they meet two conditions: first, that the element does not already exist in the set, and second, that it is bracket-independent, see Definition 16. The latter is confirmed either by checking the linear independence, as done in [Sapova2022], or by verifying that it cannot be written as an adjoint sequence. We then evaluate the rank of $\mathbf{\Gamma}_{\mathcal{A}}\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{A}|}$ for the updated set $\mathcal{A}$. The algorithm terminates if the rank is $rank(\mathbf{\Gamma}_{\mathcal{A}})=2N-4$. It is worth noting that while our code returns exactly a minimal complete pool, an MCP is often insufficient for convergence. Therefore, we typically add more starters and work with larger, Complete Pools (CP) in our simulations. Finally these CPs are used to construct the ansatz, either using fixed (NI–VQE) or iterative methods such as ADAPT-VQE or its variants (e.g batched ADAPT [Sapova2022]).
@@
All simulations mentioned above ran onto accelerated computing nodes equiped with NVIDIA A100 and H100 GPU cards. Except for H2O, all other molecules require a single 40GB card from a DGX-A100 machine to host the simulations; while H2O required 8 H100 with 80GB each, due to the storage requirements of the Hamiltonian matrix, i.e. approximately 472 Gigabytes in its Compressed Sparse Row (CSR) format. This latter molecule simulation naturally operated in parallel, and was performed on the Jean Zay supercomputer (IDRIS) at GENCI, whose recent partition is equipped with a set of 4 interconnected H1OO Cards per node, each node exchanging data through a NDR InfiniBand 400GB network.

### v2

@@ I Introduction
The evolution of quantum states is fundamentally described by unitary transformations that form a Lie group. The generators of these evolutions are elements of the corresponding Lie Algebra, which can be understood as the tangent space at the identity element of the Lie group when seen as a smooth manifold.
Lie Algebras are central in quantum computation, where any transformation can be realized as a sequence of unitary evolutions generated by operators from a specific Lie Algebra, typically $\mathfrak{su}(2^{N})$ for a $N$-qubit system. A key research area involves understanding the generation of these algebras from a minimal set of operators. For instance, when using a set of products of Pauli operators, it is known that a minimum of $2N+1$ elements in this set is required to generate the full basis of $\mathfrak{su}(2^{N})$ through nested commutation relations. The span of this basis is known as Dynamical Lie Algebra (DLA), a term used in the context of quantum optimal control [Albertini, Wiersema2024, Goh2025, fuclass, Smith2025, class2, Kazi2025], and later adopted in quantum computing. Consequently, significant past work has aimed to investigate specific properties and applications of these bases. Broadly, these efforts can be categorized into two themes: foundational classification [Wiersema2024, Goh2025, fuclass, Smith2025, class2, Kazi2025] and practical application in variational quantum algorithms [Shkolnikov2023, Sapova2022, Haidar2025].
@@ II.1 Summary of Main Lie Algebra Results
The following table summarizes different approaches from the literature for verifying MCPs, alongside our proposed method. For more details on these existing methods, we refer the reader to Supplementary Material:
@@ II.2 Preliminaries
The proof of this proposition is given in Supplementary Material.
@@ II.3 Main theorems
For sake of clarity and brevity, we will present our results without proofs. Please refer to Supplementary Material for additional details and fully worked-out proofs.
@@
The proofs for Propositions 13 and 15 are detailed in Supplementary Material. The idea behind these propositions is that we can find a sequence of contractions, represented by a matrix $\mathbf{P}$, that transforms $\mathbf{\Gamma}_{\mathcal{A}}$ into an equivalent matrix $\mathbf{\Gamma}_{equiv}$ with a known classification.
@@
The corresponding proof is given in Supplementary Material.
@@
The proofs of these theorems are given in Supplementary Material.
@@ II.4 Outlook on Applications: Quantum Chemistry on Quantum Computers
where $M\in\mathbb{N}^{*}$, and $\boldsymbol{\theta}=\left\{\theta_{i}\right\}_{i=1}^{M},~\theta_{i}\in\mathbb{R}$ is a list of parameters that completely characterize the state. $U_{i},i\in[M]$ are unitary operators, can be defined also as excitation operators acting on a reference state $|\Psi_{0}\rangle$. This algorithm’s distinct feature is its iterative approach to constructing this parametrized quantum state. At each iteration, the state is expanded by appending an exponentiated generator, which is carefully selected from a set of operator pools. The general workflow of this algorithm is described in Supplementary Material. There are many types of operator pools that can be used to construct the ADAPT-VQE ansatz, however one significant measurement overhead of ADAPT-VQE is the need to measure the energy gradient for every operator in the predefined operator pool which traditionally scales $\mathcal{O}(N^{4})$ yielding a measurement cost of $\mathcal{O}(N^{8})$. The work of [Shkolnikov2023], introduces a method to mitigate this cost by carefully constructing the operator pool, such that they use:
@@
Symmetries and properties of the pool: For molecular simulations with real-valued ansatz, the operator pool generators are restricted to odd Pauli strings (i.e odd number of Y operators in the string). The complete set of these generators forms the $\mathfrak{so}(2^{N})$ Lie Algebra, see Supplementary Material. However, molecular Hamiltonians possess additional inherent symmetries, such as the conservation of particle number, total spin, spin projection, and spatial point-group symmetries, see Supplementary Material for more details. If the pool generators do not respect these symmetries, gradient-based algorithms like ADAPT-VQE can fail. This failure manifests as vanishing gradients, symmetry roadblock [Shkolnikov2023], preventing the algorithm from starting or converging.
@@
It is worth mentioning that the set of operators that preserves all these molecular symmetries is not necessarily closed under the Lie brackets, i.e the commutator of two symmetric operators may produce a new operator that violates the desired symmetry (e.g. particle number, see Supplementary Material). Given that completeness requires the final pool to generate a mathematically closed Lie Algebra, the algebra generated from these starters will unavoidably expand to include operators that do not preserve all of the initial symmetries. We have therefore investigated which of these properties form a closed subalgebra. Our analysis indicates that the largest subalgebra that respects these constraints is defined by the two properties: the odd string requirement and the even flip property, the proof is detailed in Supplementary Material. We found numerically that the MCPs corresponding to this specific Lie subalgebra, which is a subalgebra of $\mathfrak{so}(2^{N})$, have $\mathbf{\Gamma}_{\mathcal{A}}$ matrices of rank 2N42N-4. This finding is currently numerical, as a formal proof and detailed characterization of this Lie subalgebra are beyond the scope of this work.
@@ II.5 Numerical results
These methods are performed across systems ranging from 12 to 26 qubits, from weak to strongly correlated systems. The specific molecular benchmarks include: LiH ($R=1.3\text{{\AA }}$, 12 qubits), the hydrogen chain H6 ($R=3.0\text{\AA }$, 12 qubits), H8 ($R=0.8\text{\AA }$, 16 qubits), and H2O (26 qubits, geometry defined in Supplementary Material). All simulations were performed using our Hyperion statevector emulator [hyperion]. We employed the minimal basis set (STO-3G) for most systems, while the 66-$31G$ basis set was used specifically for the H2O molecule. We assess ansatz performance across iterations using four metrics: energy accuracy, CNOT count, function evaluations (including the gradient measurements for ADAPT-VQE approaches), and the number of variational parameters.
@@
Figure 2: Convergence plots for ADAPT-VQE and MB-ADAPT-VQE methods across various molecules: LiH ($R=1.3\text{{\AA }}$, 12 qubits), the hydrogen chain H6 ($R=3.0\text{\AA }$, 12 qubits), H8 ($R=0.8\text{\AA }$, 16 qubits), and H2O (26 qubits, geometry defined in Supplementary Material). The ADAPT-VQE method uses a QEB pool of size $\mathcal{O}(N^{4})$. The composite MB-ADAPT-VQE method employs batched-ADAPT-VQE with efficient construction of MCPs of size $\mathcal{O}(N)$ for various values of $k$. Plots display energy errors versus CNOT count, function evaluations, and parameter count for different molecules.
@@
Figure 3: Convergence plots for NI-DUCC-VQE method across various molecules using different pool operators: the hydrogen chain H6 ($R=3.0\text{\AA }$, 12 qubits), H8 ($R=0.8\text{\AA }$, 16 qubits), and H2O (26 qubits, geometry defined in Supplementary Material). Plots display energy errors versus function evaluations.
@@ IV Methods
We give in Figure 5, an illustrated scheme on different steps towards a user-defined complete pool. As mentioned above, for our user-defined pool, which targets molecular Hamiltonians, the process begins with an initial set of starters. These starters must preserve specific symmetries and properties (see Supplementary Material), such as odd strings, even flip, particle number, and spin. These starters can be selected in several ways: random selection from subgroups $G\subseteq i\mathcal{P}_{N}^{*}$ that preserve desired symmetries, random selection from UCCSD operator pool ( accounting only for single and double excitations) [Sapova2022, Haidar2025], or using the UCCSD pool combined with prescreening techniques to select the most significant contributors to energy convergence [Haidar2025]. Once the starters are chosen, the complete pool is built iteratively. We add new elements that preserve the required symmetries, provided they meet two conditions: first, that the element does not already exist in the set, and second, that it is bracket-independent, see Definition 16. The latter is confirmed either by checking the linear independence, as done in [Sapova2022], or by verifying that it cannot be written as an adjoint sequence. We then evaluate the rank of $\mathbf{\Gamma}_{\mathcal{A}}\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{A}|}$ for the updated set $\mathcal{A}$. The algorithm terminates if the rank is $rank(\mathbf{\Gamma}_{\mathcal{A}})=2N-4$. It is worth noting that while our code returns exactly a minimal complete pool, an MCP is often insufficient for convergence. Therefore, we typically add more starters and work with larger, Complete Pools (CP) in our simulations. Finally these CPs are used to construct the ansatz, either using fixed (NI–VQE) or iterative methods such as ADAPT-VQE or its variants (e.g batched ADAPT [Sapova2022]).
@@
All simulations mentioned above ran onto accelerated computing nodes equipped with NVIDIA A100 and H100 GPU cards. Except for H2O, all other molecules require a single 40GB card from a DGX-A100 machine to host the simulations; while H2O required 8 H100 with 80GB each, due to the storage requirements of the Hamiltonian matrix, i.e. approximately 472 Gigabytes in its Compressed Sparse Row (CSR) format. This latter molecule simulation naturally operated in parallel, and was performed on the Jean Zay supercomputer (IDRIS) at GENCI, whose recent partition is equipped with a set of 4 interconnected H1OO Cards per node, each node exchanging data through a NDR InfiniBand 400GB network.
@@ Acknowledgments
Competing Interests
JPP is shareholder and co-founder of Qubit Pharmaceuticals. The remaining authors declare no other competing interests.